If n represents the amount of songs on MP3 player, analyzethe meaning of them expression
N +7, n-2, 4n and n÷2.

Answer by Menjax(62)

adding 7 songs to the mp3 = N+7
deleting two songs from the mp3 = N-2
4 times the amount of songs on the mp3 = 4N
half the songs on the mp3 = N/2